OpenStudy (anonymous):

Darlene bought some 30 cent and some 60 cent candy bars. Altogether she bought 10 candy bars and paid a total of $4.20. How many candy bars at each price did she buy?

OpenStudy (anonymous):

help!

OpenStudy (anonymous):

t stands for 30 cent candy bars s stands for 60 cent candy bars t + s = 10 ---> t = 10 - s .30t + .60s = 4.20 now sub 10 - s in for t in the 2nd equation .30t + .60s = 4.20 .30(10 - s) + .60s = 4.20 3 - .30s + .60s = 4.20 3 + .30s = 4.20 .30s = 4.20 - 3 .30s = 1.20 .30s/.30 = 1.20/.30 s = 4 now sub 4 in for s in the first equation... t + s = 10 t + 4 = 10 t = 10 - 4 t = 6 check... .30t + .60s = 4.20 .30(6) + .60(4) = 4.20 1.80 + 2.40 = 4.20 4.20 = 4.20 (correct) ANSWER : Darlene bought 6 thirty cent candy bars at 1.80 and she bought 4 sixty cent candy bars at 2.40 for a total of 10 candy bars costing 4.20 Do you understand how I got this answer ?